# 3D-MolT5: Towards Unified 3D Molecule-Text Modeling with 3D Molecular Tokenization

### v1

@@ 3.3 Molecular Tokenization
After obtaining the 1D SELFIES sequence $S=\{s_{i}\}_{i=1}^{m}$ and 3D sequence (i.e., the sequence of 3D tokens) $D=\{\bm{d}_{i}\}_{i=1}^{n}$, we further combine them to create the final 1D + 3D joint representation. As shown in Figure 2, for each molecule, most 1D SELFIES tokens $s_{i}$ uniquely represent an atom $a_{i}$. Each 3D token $\bm{d}_{i}$, a $k+1$ dimensional vector of nonnegative identifiers, also uniquely corresponds to an atom $a_{i}$. Thus, tokens from both 1D and 3D modalities can be aligned at the atomic level. Based on this alignment, we construct our 1D + 3D joint representation, capturing both the chemical sequence and spatial configuration of the molecule.
We define the 1D embedding $\mathbf{E}_{\text{1D}}$ for the 1D SELFIES tokens and the 3D embedding $\mathbf{E}_{\text{3D}}$ for the 3D tokens. Note that the 3D embedding $\mathbf{E}_{\text{3D}}$ is directly indexed by the 3D tokens, as each item of $\bm{d}_{i}$ is a nonnegative integer. For each molecule, each SELFIES token $s_{i}$ is mapped to its corresponding embedding vector $\mathbf{E}_{\text{1D}}(s_{i})\in\mathbb{R}^{H}$, where $H$ is the hidden dimension. For the 3D token embeddings, we first map each item $\bm{d}_{ij}$ in $\bm{d}_{i}$ to the vector $\mathbf{E}_{\text{3D}}(\bm{d}_{ij})\in\mathbb{R}^{H}$. These $k+1$ vectors $\mathbf{E}_{\text{3D}}(\bm{d}_{ij})$ are then averaged to obtain the 3D embedding for token $\bm{d}_{i}$, which is given by $\mathbf{E}_{\text{3D}}(\bm{d}_{i})=(1/k+1)\sum_{j=0}^{k}{\mathbf{E}_{\text{3D}}(\bm{d}_{ij})}$. The final joint representation $\mathbf{E}$ for each atom token is computed as:
$\mathbf{E}=\begin{cases}\mathbf{E}_{\text{1D}},&\text{if only 1D information is present}\\ \mathbf{E}_{\text{3D}},&\text{if only 3D information is present}\\ 0.5\cdot\mathbf{E_{\text{1D}}}+0.5\cdot\mathbf{E}_{\text{3D}},&\text{if both 1D and 3D information are present}\end{cases}$
This combination captures both the sequential and spatial information of the molecule, resulting in a comprehensive representation suitable for various downstream tasks.
@@ 3.4 Pre-training
With the molecular tokenization approach introduced above, we can train LMs to incorporate 3D molecular structures. Our model backbone is T5 [43], a transformer-based encoder-decoder model, which forms the foundation of our 3D-MolT5. Detailed model configurations are provided in Appendix A. For pre-training, we design two categories of tasks within a multi-task framework: (1) self-supervised denoising task to recover masked spans, and (2) translation tasks between different modalities to further enhance the model’s capability (see ablation study in Section 5). More details about pre-training configurations and datasets can be found in Appendix B.
Denoising Tasks.
Given a sequence $X=\{x_{i}\}_{i=1}^{n}$, some consecutive spans of $X$ are randomly masked by sentinel tokens, and the model learns to reconstruct these spans:
$\mathcal{L}_{\mathrm{D}}=-\sum_{t=1}^{|M|}\log P(X_{M}\mid X_{\backslash M}),$ (1)
@@
The denoising pre-training tasks are divided into two categories based on input modalities: (1) 1D denoising, including denoising on SELFIES, text, and “wrapped” text. For SELFIES, we random sample 50M molecules from the PubChem [19] database and represent them as canonical SELFIES. For text, we use both the C4 [43] dataset from the general domain and full articles from PubMed Central [5, 57] in the biomedical domain. The concept of “wrapped” text is derived from MolXPT [29], where the molecules mentioned in the text are replaced with SELFIES. As in  [29, 39, 40], training on such “wrapped” text is helpful as the context is full of descriptions about the molecule. We follow the same pipeline as MolXPT to detect molecules mentioned in text using BERN2 [50] on the PubMed abstracts, and then suffix them with corresponding SELFIES. (2) 1D + 3D joint denoising, which involves denoising on summed 1D SELFIES tokens and 3D tokens, with the goal of recovering the 1D SELFIES tokens [^2] In our work, since the 3D molecular token for each atom contains $k+1$ different items, we do not aim to predict the 3D molecular tokens for the denoising tasks, which is same for the translation tasks.. We use the PCQM4Mv2 dataset from the OGB Large Scale Challenge [17] for this task, which includes 3.37M DFT-calculated [16] 3D molecular structures that are commonly regarded as ground-state 3D structures.
Translation Tasks.
In addition to denoising tasks, we also add translation tasks between modalities to enhance the representation learning,
$\mathcal{L}_{\mathrm{T}}=-\sum_{t=1}^{|Y|}\log P(Y\mid X),$ (2)
@@
Concretely, we incorporate three translation tasks from $X$ to $Y$ to further connect different modalities: (1) 3D to 1D translation. We use the same PCQM4Mv2 dataset as in 1D + 3D denoising in Section 3.4. The input is the sequence of 3D token representations $\mathbf{E}_{\text{3D}}$ of the molecule, and the output is the corresponding 1D SELFIES. (2) 3D molecule to text translation. We use the pre-training split of the PubChem dataset collected by 3D-MoLM [25], which includes 298K 3D molecule-text pairs from the PubChem [19] database. The input is the summed 1D and 3D tokens of the molecule, and the output is the corresponding textual description. (3) Text to 1D molecule translation. We use the same data as (2), but the input is the textual description of the molecule, and the output is the corresponding 1D SELFIES.
@@ 4 Experiments
We evaluate 3D-MolT5 on three types of text-based molecule-related downstream tasks: (1) molecular property prediction, including both 3D-independent properties (e.g., molecular weight, LogP) and 3D-dependent properties (e.g., HOMO-LUMO gap); (2) 3D molecule captioning; (3) text-based molecule generation. All downstream data is formatted as instructions, with tasks framed as conditional text or molecule generation based on the input instructions. The training objective is the same as the standard cross-entropy loss used in pre-training the language model.
Following 3D-MoLM [25] and Mol-Instructions [14], we present performances of two variants of 3D-MolT5: Specialist, which is fine-tuned only on a specific task; Generalist, which is fine-tuned in a multi-task setup. We maintain the same multi-task setup as the baselines for a fair comparison. More details about the downstream datasets, Generalist settings, and baselines are in Appendix C.
@@ 4.1 Molecular Property Prediction
Following 3D-MoLM [25], we assess 3D-MolT5 on two types of molecular property prediction tasks. (1) For computed property prediction, we focus on the MAE performance by extracting the predicted numerical value of the property from the output text. (2) For descriptive property prediction, we evaluate text similarity metrics between the predicted text and the ground truth.
@@ 4.1.1 Computed Property Prediction
Setup. We adopt three datasets to evaluate the performance of 3D-MolT5 on computed property prediction task, including QM9 [47, 14], PubChemQC [34, 59], and PubChem [19]. The QM9 [47, 14] dataset includes over 130,000 molecules with their ground-state 3D structures obtained through DFT computations [16], each molecule having fewer than nine heavy atoms. It is widely used for quantum property prediction, including HOMO, LUMO, and HOMO-LUMO gap (H-L gap). The PubChemQC [34, 59] dataset is on a larger scale, containing 3.37M molecules with more heavy atoms with their DFT-calculated [16] 3D structures. We use the same quantum properties as QM9 for PubChemQC. The PubChem [19] database also provides various molecular properties, and we follow [25] to select four properties that can be inferred from 1D or 2D molecular information.
@@
For the QM9 dataset, we use instruction data from Mol-Instructions [14]. For the PubChemQC and PubChem datasets, we use instruction data constructed by 3D-MoLM [25].
@@
Baselines & Evaluation. We compare 3D-MolT5 against three types of baseline models based on input modalities. For 1D sequence-based models, we include Llama2-7B [52], Vicuna [8], Mol-Instructions [8], and BioT5+ [40]. For 2D graph-based models, we include 2D-MoLM [25] and InstructMol [6]. For the 3D structure-based model, we compare with Uni-Mol [63] and 3D-MoLM [25]. Note that models based on 2D graphs or 3D structures can also concurrently accept 1D sequences as input [25, 6].
@@
Results. The computed property results for the PubChem and PubChemQC datasets are shown in Table 1, and for QM9 in Table 2. Key observations from the results include: (1) 3D-MolT5 achieves superior performances on 3D-independent properties over all baseline methods on the PubChem [19] dataset. For molecular hydrophobicity (LogP), determined by 1D and 2D features like functional groups, molecular connectivity, and topology, 3D-MolT5 consistently surpasses the LMs trained on 1D, 2D, and 3D molecular information. (2) 3D-MolT5 exhibits significant improvements on 3D-dependent properties. For energy properties including HOMO, LUMO, and HOMO-LUMO gap, which are largely determined by 3D molecular structure, 3D-MolT5 shows strong performance enhancements. For the Specialist version on the PubChemQC [34] dataset, the improvements over the previous SOTA methods [63, 25] are 0.18 eV, 0.17 eV, and 0.13 eV, respectively. For the Generalist version on the QM9 [47] dataset, the average improvement is 0.0008 Ha. (3) Compared to Uni-Mol [63], 3D-MolT5 also shows consistent improvement. Uni-Mol [63] is specially designed for 3D molecular representation learning and pre-trained on large-scale (209M) 3D molecular data. The superiority of 3D-MolT5 illustrates the benefit of unified 3D molecule-text modeling. Through comprehensive pre-training, 3D-MolT5 gains both structural knowledge from molecules and contextual knowledge from the biological literature, enhancing its generalization to molecular property prediction tasks. (4) Compared to 3D-MoLM [25], 3D-MolT5 also exhibits continuous enhancement. 3D-MoLM [25] integrates the molecular encoder Uni-Mol [63] with Llama2-7B [52] as the language decoder for 3D molecule-text interpretation. With a 3D molecule-to-text projector, the molecular encoder and LM can be aligned by training on molecule-text pairs. The superiority of 3D-MolT5 can be attributed to multi-task pre-training and 3D molecular tokenization, which significantly improve 3D-MolT5’s ability to handle complex fine-grained 3D molecular structures and enhance cross-modal understanding. (5) On PubChemQC and PubChem datasets, the Generalist version of 3D-MolT5 also outperforms all baselines but slightly underperforms compared to the Specialist version, likely due to the nature of multi-task training. These results demonstrate the capability of 3D-MolT5 to handle multiple tasks simultaneously as a Generalist.
@@
Table 2: MAE results on computed property prediction tasks on QM9 [47] dataset. * means direct inference without further fine-tuning.
Model HOMO (Ha) LUMO (Ha) H-L Gap (Ha) Avg $\downarrow$
@@ 4.1.2 Descriptive Property Prediction
Setup. For the descriptive property prediction task, we evaluate the performance of our proposed 3D-MolT5 on the PubChem dataset, as constructed by 3D-MoLM [25]. Unlike computed property prediction, this task entails generating descriptions of molecular 1D, 2D, and 3D properties in natural language, necessitating accurate and contextually relevant textual output.
Baselines & Evaluation. We compare 3D-MolT5 against the same baselines evaluated on PubChemQC [34] and PubChem [19] datasets as Section 4.1.1. Following MolT5 [13], the evaluation metrics are commonly used text generation metrics, including BLEU [38], ROUGE [26], and METEOR [3], to measure the similarity between the generated property descriptions and ground truth.
Results. The results for the PubChem dataset are presented in Table 3. From the table, we can have several findings: (1) 3D-MolT5 outperforms all baseline methods. Specifically, the Specialist version shows significant improvements, with BLEU-2 and ROUGE-L scores increasing by 19.24 and 17.84, respectively. The Generalist version also shows substantial gains, with improvements of 18.03 and 15.4 in BLEU-2 and ROUGE-L, respectively. (2) The Generalist version of 3D-MolT5 outperforms all baselines but slightly underperforms compared to the Specialist version.
@@
Setup. We use PubChem [25] dataset for 3D molecule captioning to evaluate 3D-MolT5’s proficiency to understand the 3D molecular structure. This dataset contains approximately 15,000 3D molecular structure-text pairs source from the PubChem database [19]. Each input molecule has an optimized 3D structure, and the output is the corresponding caption that may include 3D-related descriptions. Specifically, the molecular captions include both molecular names and descriptions to assess the model’s capability in name prediction [15] and description prediction [13], offering more comprehensive evaluation compared to the descriptive properties in Section 4.1.2.
Baselines & Evaluation. The compared baselines are classified into three types based on input modalities. For 1D sequence-based models, we include MolT5 [13] and Llama2-7B [52]. For 2D graph-based models, we include MoMu [49] and 2D-MoLM [25]. For 3D structure-based models, we compare with 3D-MoLM [25]. Note that the 2D graph or 3D structure-based models may also take a 1D sequence as input simultaneously. The evaluation metrics are the same as those in Section 4.1.2
Results. Table 4 shows the performances for the 3D molecule captioning task. 3D-MolT5 demonstrates excellent results, surpassing all baselines. Compared to 3D-MoLM [25], 3D-MolT5 achieves an improvement of approximately 11 points in ROUGE-L and METEOR scores for both the Specialist and Generalist versions. 3D-MolT5 also exceeds baselines with 1D and 2D information, highlighting the importance of 3D structure for molecule understanding and captioning. This further validates the efficacy of our pre-training on 1D SELFIES, 3D structure, and text with 3D molecular tokenization.

### v2

@@ 3.3 Molecular Tokenization
After obtaining the 1D SELFIES sequence $S=\{s_{i}\}_{i=0}^{m-1}$ and the sequence of 3D tokens $\mathbf{D}=\{\bm{d}_{i}\}_{i=0}^{n-1}$, we combine them to form the final 1D + 3D joint representation. As depicted in Figure 2, for each molecule, most 1D SELFIES tokens $s_{i}$ uniquely represent an atom $a_{i}$. Similarly, each 3D token $\bm{d}_{i}$, a $k+1$ dimensional vector of non-negative identifiers, also uniquely corresponds to an atom $a_{i}$. Thus, tokens from both 1D and 3D modalities can be aligned at the atomic level. Based on this alignment, we construct the 1D + 3D joint representation, capturing both the chemical sequence and the spatial configuration of the molecule.
We define the 1D embedding $\mathbf{E}_{\text{1D}}$ for the 1D SELFIES tokens and the 3D embedding $\mathbf{E}_{\text{3D}}$ for the 3D tokens. The 3D embedding $\mathbf{E}_{\text{3D}}$ is directly indexed by the 3D tokens, as each component of $\bm{d}_{i}$ is a non-negative integer. For each molecule, each SELFIES token $s_{i}$ is mapped to its corresponding embedding vector $\mathbf{E}_{\text{1D}}(s_{i})\in\mathbb{R}^{H}$, where $H$ denotes the hidden dimension. For the 3D token embeddings, we map each component $\bm{d}_{i,j}$ in $\bm{d}_{i}$ to the vector $\mathbf{E}_{\text{3D}}(\bm{d}_{i,j})\in\mathbb{R}^{H}$. These $k+1$ vectors $\mathbf{E}_{\text{3D}}(\bm{d}_{i,j})$ are then averaged to compute the 3D embedding for token $\bm{d}_{i}$, given by $\mathbf{E}_{\text{3D}}(\bm{d}_{i})=(1/k+1)\sum_{j=0}^{k}{\mathbf{E}_{\text{3D}}(\bm{d}_{i,j})}$. The final joint representation $\mathbf{E}$ for each token is determined based on the available information: if only 1D information is present, $\mathbf{E}=\mathbf{E}_{\text{1D}}$; if only 3D information is available, $\mathbf{E}=\mathbf{E}_{\text{3D}}$; and if both 1D and 3D information are present, then $\mathbf{E}=\frac{1}{2}\mathbf{E}_{\text{1D}}+\frac{1}{2}\mathbf{E}_{\text{3D}}$. This combination captures both the sequential and spatial information of the molecule, producing a comprehensive representation suitable for various downstream tasks.
@@ 3.4 Pre-training
Using the molecular tokenization approach described above, now we can train LMs on text sequences, molecular sequences, and also molecular structures all in tokens. Our LM backbone is T5 (Raffel et al., 2020), a transformer-based encoder-decoder architecture, which serves as the foundation for 3D-MolT5. Detailed model configurations are provided in Appendix D. For pre-training, we design two categories of tasks within a multi-task framework: (1) self-supervised denoising tasks aimed at recovering masked spans, and (2) translation tasks between different modalities to further enhance the model’s capability (see ablation study in Section 5). Additional details regarding pre-training, including loss functions, configurations, and datasets, are provided in Appendix E.
@@
Denoising Tasks. The denoising pre-training tasks are divided into two categories based on input modalities: (1) 1D denoising, which includes denoising on SELFIES, text, and “wrapped” text. For SELFIES, we random sample 50M molecules from the PubChem (Kim et al., 2019) database and represent them as canonical SELFIES. For text, we use both the C4 (Raffel et al., 2020) dataset from the general domain and full articles from PubMed Central (Canese & Weis, 2013; White, 2020) in the biomedical domain. The concept of “wrapped” text is adapted from MolXPT (Liu et al., 2023b), where molecules mentioned in the text are replaced with their SELFIES. As demonstrated in  Liu et al. (2023b); Pei et al. (2023; 2024a), training on such “wrapped” text is beneficial, as the context is rich with molecular descriptions. We follow the same pipeline as MolXPT to detect molecules mentioned in the text using BERN2 (Sung et al., 2022) on the PubMed abstracts, appending them with their corresponding SELFIES. (2) 1D + 3D joint denoising, which involves denoising on the combined 1D SELFIES tokens and 3D tokens, aiming at recovering the 1D SELFIES tokens [^2] In our work, since each 3D molecular token for an atom contains $k+1$ different components, we do not attempt to predict the 3D molecular tokens for denoising tasks, as is also the case for the translation tasks.. We use the PCQM4Mv2 dataset from the OGB Large Scale Challenge (Hu et al., 2021) for this task, which includes 3.37M DFT-calculated (Geerlings et al., 2003) 3D molecular structures.
@@
Translation Tasks. We incorporate three translation tasks simultaneously to further bridge different modalities: (1) 3D to 1D translation. We use the same PCQM4Mv2 dataset as in the 1D + 3D denoising. The input is the sequence of 3D token representations $\mathbf{E}_{\text{3D}}$ of the molecule, and the output is the corresponding 1D SELFIES. (2) 3D molecule to text translation. We use the pre-training split of the PubChem dataset collected by 3D-MoLM (Li et al., 2023c), which contains 298K 3D molecule-text pairs from the PubChem (Kim et al., 2019) database. The input is the combined 1D and 3D tokens of the molecule, and the output is the corresponding textual description. (3) Text to 1D molecule translation. We use the same data as (2), but the input is the textual description of the molecule, and the output is the corresponding 1D SELFIES.
@@ 4 Experiments
We evaluate 3D-MolT5 on three types of text-based molecule-related downstream tasks: (1) molecular property prediction, including both 3D-independent properties (e.g., molecular weight, LogP) and 3D-dependent properties (e.g., HOMO-LUMO gap); (2) 3D molecule captioning; (3) text-based molecule generation. In Appendix B, we further validate the effectiveness of 3D-MolT5 across additional tasks and benchmarks. All downstream data is formatted as instructions, with tasks framed as conditional text or molecule generation based on the input instructions. The training objective remains the standard cross-entropy loss, consistent with pre-training.
Following 3D-MoLM (Li et al., 2023c) and Mol-Instructions (Fang et al., 2023), we present results for two variants of 3D-MolT5: Specialist, fine-tuned for a specific task; and Generalist, fine-tuned in a multi-task setup. To ensure fair comparisons, we use the same multi-task setup as the baseline models. More details on the downstream datasets, Generalist settings, and baseline methods are provided in Appendix F.
@@ 4.1 Molecular Property Prediction
Following 3D-MoLM (Li et al., 2023c), we assess 3D-MolT5 on two types of molecular property prediction tasks: (1) Computed property prediction: We focus on the MAE performance by extracting the predicted numerical value of the property from the generated text. (2) Descriptive property prediction: We evaluate text similarity metrics between the predicted text and the ground truth.
@@ 4.1.1 Computed Property Prediction
Setup. We use three datasets to evaluate the performance of 3D-MolT5 on computed property prediction task: QM9 (Ruddigkeit et al., 2012; Fang et al., 2023), PubChemQC (Maho, 2015; Xu et al., 2021), and PubChem (Kim et al., 2019). The QM9 (Ruddigkeit et al., 2012; Fang et al., 2023) dataset contains over 130,000 molecules with ground-state 3D structures obtained through DFT computations (Geerlings et al., 2003), each molecule having fewer than nine heavy atoms. This dataset is widely used for quantum property prediction, including HOMO, LUMO, and HOMO-LUMO gap (H-L gap). The PubChemQC (Maho, 2015; Xu et al., 2021) dataset is larger in scale, containing 3.37M molecules with more heavy atoms, along with their DFT-calculated (Geerlings et al., 2003) 3D structures. We use the same quantum properties as QM9 for PubChemQC. The PubChem (Kim et al., 2019) database also provides various molecular properties. Following 3D-MoLM (Li et al., 2023c), we select four properties that can be inferred from 1D or 2D molecular information.
@@
For the QM9 dataset, we use instruction data from Mol-Instructions (Fang et al., 2023). For PubChemQC and PubChem datasets, we use instruction data constructed by 3D-MoLM (Li et al., 2023c).
@@
Baselines & Evaluation. We compare 3D-MolT5 against three types of baseline models, categorized by input modalities. For 1D sequence-based models, we include Llama2-7B (Touvron et al., 2023), Vicuna (Chiang et al., 2023), Mol-Instructions (Chiang et al., 2023), and BioT5+ (Pei et al., 2024a). For 2D graph-based models, we incorporate 2D-MoLM (Li et al., 2023c) and InstructMol (Cao et al., 2023). For the 3D structure-based model, we compare against Uni-Mol (Zhou et al., 2023) and 3D-MoLM (Li et al., 2023c). Note that models based on 2D graphs or 3D structures can also accept 1D sequences as input (Li et al., 2023c; Cao et al., 2023).
@@
Results. The computed property results for the PubChem and PubChemQC datasets are presented in Table 1, and for QM9 in Table 2. Key findings from the results include: (1) 3D-MolT5 outperforms all baseline methods on 3D-independent properties in the PubChem dataset (Kim et al., 2019). For molecular hydrophobicity (LogP), which depends on 1D and 2D features such as functional groups, molecular connectivity, and topology, 3D-MolT5 consistently surpasses the LMs trained on 1D, 2D, and 3D molecular information. (2) 3D-MolT5 exhibits substantial improvements in predicting 3D-dependent properties. For energy properties including HOMO, LUMO, and HOMO-LUMO gap, which are primarily determined by 3D molecular structures, 3D-MolT5 shows significant performance enhancements. For the Specialist version on the PubChemQC (Maho, 2015) dataset, the improvements over the previous SOTA methods (Zhou et al., 2023; Li et al., 2023c) are 0.18 eV, 0.17 eV, and 0.13 eV, respectively. For the Generalist version on the QM9 (Ruddigkeit et al., 2012) dataset, the average improvement is 0.0008 Ha. (3) Compared to Uni-Mol (Zhou et al., 2023), 3D-MolT5 exhibits consistent improvements. Uni-Mol (Zhou et al., 2023) is specially designed for 3D molecular representation learning and pre-trained on large-scale (209M) 3D molecular data. The superiority of 3D-MolT5 demonstrates the benefit of unified 3D molecule-text modeling. By integrating structural knowledge from molecules with contextual knowledge from biological literature through comprehensive pre-training, 3D-MolT5 enhances its generalization to molecular property prediction tasks. (4) 3D-MolT5 also continuously improves upon 3D-MoLM (Li et al., 2023c), which integrates the Uni-Mol (Zhou et al., 2023) molecular encoder with Llama2-7B (Touvron et al., 2023) as the language decoder through a projector for 3D molecule-text interpretation. The superiority of 3D-MolT5 can be attributed to the extensive interaction between 3D molecular structure and text during multi-task pre-training and our joint tokenization, which significantly improves 3D-MolT5’s ability to handle complex fine-grained 3D molecular structures and enhance cross-modal understanding. (5) On PubChemQC and PubChem datasets, the Generalist version of 3D-MolT5 also outperforms all baselines, though it slightly underperforms compared to the Specialist, likely due to task conflicts in multi-task training. These results demonstrate 3D-MolT5’s ability to handle multiple tasks concurrently as a Generalist.
@@
Table 2: MAE results on computed property prediction tasks on QM9 (Ruddigkeit et al., 2012) dataset. * means direct inference without further fine-tuning.
Model HOMO (Ha) LUMO (Ha) H-L Gap (Ha) Avg (Ha)
@@ 4.1.2 Descriptive Property Prediction
Setup. For the descriptive property prediction task, we evaluate the performance of 3D-MolT5 on the PubChem dataset, as constructed by 3D-MoLM (Li et al., 2023c). Unlike computed property prediction, this task involves generating natural language descriptions of molecular 1D, 2D, and 3D properties, necessitating accurate and contextually relevant textual output.
Baselines & Evaluation. We compare 3D-MolT5 against the same baselines used for the PubChemQC (Maho, 2015) and PubChem (Kim et al., 2019) datasets as described in Section 4.1.1. Following MolT5 (Edwards et al., 2022), we employ widely used text generation metrics, including BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and METEOR (Banerjee & Lavie, 2005), to assess the similarity between the generated property descriptions and the ground truth.
Results. The results for the PubChem dataset are shown in Table 3. From the table, we have several observations: (1) 3D-MolT5 outperforms all baseline methods. The Specialist version shows significant improvements, with BLEU-2 and ROUGE-L scores increasing by 19.24 and 17.84, respectively. The Generalist version also shows substantial gains, with BLEU-2 and ROUGE-L improvements of 18.03 and 15.4 points, respectively. (2) The Generalist version of 3D-MolT5 surpasses all baselines but slightly underperforms compared to the Specialist version.
@@
Setup. We use PubChem (Li et al., 2023c) dataset for 3D molecule captioning to evaluate 3D-MolT5’s ability to understand the 3D molecular structure. This dataset contains approximately 15,000 3D molecular structure-text pairs sourced from the PubChem database (Kim et al., 2019). Specifically, the molecular captions include both molecular names and 3D-related descriptions to assess the model’s capability in name prediction (Favre & Powell, 2013) and description prediction (Edwards et al., 2022), offering a more comprehensive evaluation compared to the descriptive properties.
Baselines & Evaluation. The compared baselines are classified into three categories based on input modalities. For 1D sequence-based models, we include MolT5 (Edwards et al., 2022) and Llama2-7B (Touvron et al., 2023). For 2D graph-based models, we compare against MoMu (Su et al., 2022), 2D-MoLM (Li et al., 2023c), UniMoT (Zhang et al., 2024a), and MolX (Le et al., 2024). For 3D structure-based models, we incorporate 3D-MoLM (Li et al., 2023c). Note that the 2D graph or 3D structure-based models may also take a 1D sequence as input simultaneously. The evaluation metrics remain consistent with those in Section 4.1.2.
Results. Table 4 shows the results for the 3D molecule captioning task. 3D-MolT5 demonstrates superior results, surpassing all baseline methods. Compared to 3D-MoLM (Li et al., 2023c), 3D-MolT5 achieves an improvement of approximately 11 points in ROUGE-L and METEOR scores for both the Specialist and Generalist versions. Furthermore, 3D-MolT5 also exceeds baselines with 1D and 2D information, highlighting the importance of 3D structure information for molecule understanding and captioning. This further validates the efficacy of our unified pre-training on 1D SELFIES, 3D structure, and text with 3D molecular tokenization.